Spec’s, a distributor for S. Pellegrino Corporation, sells S. Pellegrino mineral water to grocery stores around the world. The cost of each bottle depends on when Spec’s purchases the water. See the cost structure below:

Mar. Purchases 10 cases @ $30
Jul. Purchases 20 cases @ $40
Aug Purchases 30 cases @ $50

Spec’s sells 35 cases in September.

What is the gross margin under the FIFO method if each case is sold for $75?
Answer #1

Cost of Goods Sold under fifo

= (10*30) + (20*40) + (5*50)

= 1350

Sales = 35*75 = 2625

Gross margin = Sales - Cost of Goods Sold

= 2625-1350

= 1275